# Topologically protected negative entanglement

### v3

@@ I.1.2 Effect of bulk states on entanglement scaling due to topological EPs
For the general form of the Hamiltonian $H(k,k_{y})$ as defined in Eq. (S3), when considering OBCs in the $y$-direction with size $L_{y}$, we obtain:
$H_{\text{y-OBC}}(k)=\begin{pmatrix}h_{0}^{y}&h_{+}^{y}&\cdots&0&0\\ h_{-}^{y}&h_{0}^{y}&h_{+}^{y}&\cdots&0\\ 0&h_{-}^{y}&h_{0}^{y}&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&0&\cdots&h_{-}^{y}&h_{0}^{y}\end{pmatrix}_{L_{y}\times L_{y}},\text{with }h_{+}^{y}=\begin{pmatrix}0&0&0&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&0&0&0\end{pmatrix},\quad h_{-}^{y}=\begin{pmatrix}0&0&1&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&1&0&0\end{pmatrix},\text{ and}\quad h_{0}^{y}(k)=\begin{pmatrix}0&R\\ L&0\end{pmatrix},$ (S4)
@@
$R=\begin{pmatrix}f+\delta+Z&2Z-i(Z+\sin(k+k_{0})\\ i(Z-\sin(k+k_{0}))&f+\delta-Z\end{pmatrix},\quad L=\begin{pmatrix}f-\delta+Z&-i(Z-\sin(k+k_{0})\\ 2Z+i(Z+\sin(k+k_{0}))&f-\delta-Z\end{pmatrix},$ (S5)
and $f=\cos(k+k_{0})+\cos k_{0}-M$. For this $H_{\text{y-OBC}}(k)$, the topological edge bands exhibit two different types of dispersion relations around the EP: (1) a linear dispersion $E_{e}(\delta k)\sim\delta k$, as shown in Fig. 1(c), and (2) square-root dispersion $E_{e}(\delta k)\sim\sqrt{\delta k}$, as depicted in Fig. S1(a).
To explain why the dispersion can exhibit two qualitatively different behavior (i.e. becomes linear when we set $Z=1,\delta=M-Z$ in Eq. (S3)), we expand the momentum $k$ around the EP as: $k\rightarrow 0+\delta k$. We shall prove that:
$\det[H_{\text{y-OBC}}(\delta k)]\sim\begin{cases}\delta k^{2},&\text{with }Z=1,\delta=M-Z\\ \delta k,&\text{other parameters}\end{cases}.$ (S6)
This is connected to the dispersion of the topological bands because the latter are the only ones that vanish (are gapless). To elaborate, we know that the determinant is equal to the product of all eigenenergies as:
$\det[H_{\text{y-OBC}}(k)]=E_{e_{1}}(k)E_{e_{2}}(k)\prod_{n}E_{n}(k),$ (S7)
@@
To derive Eq. (S6), we first expand $H_{\text{y-OBC}}$ around $k=0$. Since $k$ appears only in the diagonal element $h_{0}^{y}$, we need to expand the $R$ and $L$ matrices [Eq. (S5)]. By substituting $\sin k_{0}=Z$ into Eq. (S5), we obtain
$\sin(k_{0}+\delta k)\approx Z+\cos k_{0}\delta k-\frac{Z}{2}\delta k^{2},\quad\quad\cos(k_{0}+\delta k)\approx\cos k_{0}-Z\delta k-\frac{\cos k_{0}}{2}\delta k^{2},$ (S8)
@@
$R(\delta k)=\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}=\begin{pmatrix}f_{0}+\delta+Z&2Z(1-i)\\ 0&f_{0}+\delta-Z\end{pmatrix}+\begin{pmatrix}-Z&-i\cos k_{0}\\ -i\cos k_{0}&-Z\end{pmatrix}\delta k+\begin{pmatrix}-\frac{1}{2}\cos k_{0}&i\frac{Z}{2}\\ i\frac{Z}{2}&-\frac{1}{2}\cos k_{0}\end{pmatrix}\delta k^{2}+\mathcal{O}(\delta k^{3}),$ (S9)
@@
$L(\delta k)=\begin{pmatrix}L_{11}&L_{12}\\ L_{21}&L_{22}\end{pmatrix}=\begin{pmatrix}f_{0}-\delta+Z&0\\ 2Z(1+i)&f_{0}-\delta-Z\end{pmatrix}+\begin{pmatrix}-Z&i\cos k_{0}\\ i\cos k_{0}&-Z\end{pmatrix}\delta k+\begin{pmatrix}-\frac{1}{2}\cos k_{0}&-i\frac{Z}{2}\\ -i\frac{Z}{2}&-\frac{1}{2}\cos k_{0}\end{pmatrix}\delta k^{2}+\mathcal{O}(\delta k^{3}).$ (S10)
@@
$\det\begin{pmatrix}h_{0}&h_{+}&\cdots&0\\ h_{-}&h_{0}&h_{+}&\vdots\\ 0&h_{-}&\ddots&h_{+}\\ 0&\cdots&h_{-}&h_{0}\end{pmatrix}_{L_{y}}$ $=$ $\det(h_{0})\det\begin{pmatrix}h_{0}&h_{+}&\cdots&0\\ h_{-}&h_{0}&h_{+}&\vdots\\ 0&h_{-}&\ddots&h_{+}\\ 0&\cdots&h_{-}&\color[rgb]{0,0,1}h_{0}^{(1)}\end{pmatrix}_{L_{y}-1}=\det(h_{0})\det(h_{0}^{(1)})\det\begin{pmatrix}h_{0}&h_{+}&\cdots&0\\ h_{-}&h_{0}&h_{+}&\vdots\\ 0&h_{-}&\ddots&h_{+}\\ 0&\cdots&h_{-}&\color[rgb]{0,0,1}h_{0}^{(2)}\end{pmatrix}_{L_{y}-2}$ (S11)
@@
in which
$h_{0}^{(1)}$ $=$ $h_{0}-h_{+}h_{0}^{-1}h_{-}=\begin{pmatrix}0&R^{(1)}\\ L^{(1)}&0\end{pmatrix},\quad\quad R^{(1)}=R-\begin{pmatrix}0&0\\ (R^{-1})_{21}&0\end{pmatrix},\quad L^{(1)}=L-\begin{pmatrix}0&(L^{-1})_{12}\\ 0&0\end{pmatrix}$
$h_{0}^{(2)}$ $=$ $h_{0}-h_{+}[h_{0}^{(1)}]^{-1}h_{-}=\begin{pmatrix}0&R^{(2)}\\ L^{(2)}&0\end{pmatrix},\quad R^{(2)}=R-\begin{pmatrix}0&0\\ [(R^{(1)})^{-1}]_{21}&0\end{pmatrix},\quad L^{(2)}=L-\begin{pmatrix}0&[(L^{(1)})^{-1}]_{12}\\ 0&0\end{pmatrix},$ (S12)
and so forth $h_{0}^{(n)}=h_{0}-h_{+}[h_{0}^{(n-1)}]^{-1}h_{-}$. Now, to calculate Eq. (S11), we need to determine how $\text{det}(h_{0}^{(n)})$ varies with $\delta k$.
@@
    $R(\delta k)=\begin{pmatrix}-\delta k&2(1-i)+\frac{i}{2}\delta k^{2}\\ \frac{i}{2}\delta k^{2}&-2-\delta k\end{pmatrix}+\mathcal{O}(\delta k^{3}),\quad L(\delta k)=\begin{pmatrix}2(1-M)-\delta k&-\frac{i}{2}\delta k^{2}\\ 2(1+i)-\frac{i}{2}\delta k^{2}&-2M\end{pmatrix}+\mathcal{O}(\delta k^{3}),$ (S13)
@@
    $\det(h_{0})=\det(R)\det(L)\approx-8M(1-M)\color[rgb]{0,0,1}\delta k.$ (S14)
    Importantly, the fact that $\det{R}$ is proportional to $2\delta k$ and vanishes as $\delta k\rightarrow 0$ will turn out to be crucial. According to Eq. (S12), for the first iteration, $\det(h_{0}^{(1)})$, can be calculated using $R^{(1)}$ and $L^{(1)}$
    $R^{(1)}$ $=$ $\begin{pmatrix}R_{11}&R_{12}\\ R_{21}-(R^{-1})_{21}&R_{22}\end{pmatrix}=\begin{pmatrix}R_{11}&R_{12}\\ (1+\frac{1}{\det(R)})R_{21}&R_{22}\end{pmatrix}\approx\begin{pmatrix}-\delta k&2(1-i)+\frac{i}{2}\delta k^{2}\\ \frac{i}{4}\delta k+\frac{i}{2}\delta k^{2}&-2-\delta k\end{pmatrix}+\mathcal{O}(\delta k^{3}),$
    $L^{(1)}$ $=$ $\begin{pmatrix}L_{11}&L_{12}-(L^{-1})_{12}\\ L_{21}&L_{22}\end{pmatrix}=\begin{pmatrix}L_{11}&\left(1-\frac{1}{4M(1-M)}\right)L_{12}\\ L_{21}&L_{22}\end{pmatrix}\approx\begin{pmatrix}2(1-M)-\delta k&-\frac{i}{2}\left(1-\frac{1}{4M(1-M)}\right)\delta k^{2}\\ 2(1+i)-\frac{i}{2}\delta k^{2}&-2M\end{pmatrix}+\mathcal{O}(\delta k^{3}),$ (S15)
@@
    $𝛿𝑘\det(h_{0}^{(1)})=\det(R^{(1)})\det(L^{(1)})\approx-4M(1-M)\left(2+\frac{1+i}{2}\right)\color[rgb]{0,0,1}{\delta k}.$ (S16)
@@
    $[R^{(2)}]_{21}$ $=$ $R_{21}-[(R^{(1)})^{-1}]_{21}=R_{21}+\frac{R^{(1)}_{21}}{\det(R^{(1)})}\approx\frac{i}{10+2i}=\text{Const},\quad\det(R^{(2)})\approx-\frac{1+i}{5+i}=\text{Const},$
    $[R^{(n)}]_{21}$ $=$ $R_{21}-[(R^{(n-1)})^{-1}]_{21}=R_{21}+\frac{R^{(n-1)}_{21}}{\det(R^{(n-1)})}=\frac{i}{2}\delta k^{2}+\frac{\text{Const}}{\text{Const}}\approx\text{Const},$
    $\det(R^{(n)})|_{n\geq 2}$$\approx$ $2(1-i)*[R^{(n)}]_{21}=\text{Const},$ (S17)
@@
    $[L^{(2)}]_{12}$ $=$ $L_{12}-[(L^{(1)})^{-1}]_{12}=-\frac{i\delta k^{2}}{2}+\frac{L_{12}^{(1)}}{\det(L^{(1)})}\approx\frac{i\delta k^{2}}{2}\left(\frac{1-\frac{1}{4M(1-M)}}{4M(1-M)}-1\right),\det(L^{(2)})\approx-4M(1-M)=\text{Const},$
    $[L^{(n)}]_{12}$ $=$ $L_{12}+\frac{L_{12}^{(n-1)}}{\det(L^{(n-1)})}\sim\delta k^{2},\quad\quad\det(L^{(n)})|_{n\geq 2}\approx-2(1-M)*2M=-4M(1-M)=\text{Const}.$ (S18)
@@
    $\det(H_{\text{y-OBC}})=\det(h_{0})\det(h_{0}^{(1)})\cdots\det(h_{0}^{(L_{y}-1)})\sim{\color[rgb]{0,0,1}\delta k^{2}}.$ (S19)
  • For most other generic parameter combinations, such as the case shown in Fig. S1(a), the constant term in $\det(R)$ does not vanish i.e., $\det(R)\rightarrow\text{Const}$ as $\delta k\rightarrow 0$. Below we show how that leads to a qualitatively different dispersion. We have
    $\det(R)$$\approx$ $(f_{0}+\delta)^{2}-Z^{2}-2Z\big{(}f_{0}+\delta-(1+i)\cos k_{0}\big{)}\delta k=C_{0}^{R}+C_{1}^{R}\delta k,$
    $\det(L)$$\approx$ $(f_{0}-\delta)^{2}-Z^{2}-2Z\big{(}f_{0}-\delta+(1-i)\cos k_{0}\big{)}\delta k=C_{0}^{L}+C_{1}^{L}\delta k,$
    $\det(h_{0})$ $=$ $\det(R)\det(L)\approx C_{0}^{R}C_{0}^{L}+(C_{0}^{R}C_{1}^{L}+C_{0}^{L}C_{1}^{R})\color[rgb]{0,0,1}\delta k.$ (S20)
@@
    $R^{(1)}$ $=$ $\begin{pmatrix}R_{11}&R_{12}\\ R_{21}-(R^{-1})_{21}&R_{22}\end{pmatrix}\approx\begin{pmatrix}R_{11}&R_{12}\\ (1+\frac{1}{C_{0}^{R}})R_{21}&R_{22}\end{pmatrix},\quad\det(R^{(1)})\approx C_{0}^{(1)R}+C_{1}^{(1)R}\delta k,$
    $L^{(1)}$ $=$ $\begin{pmatrix}L_{11}&L_{12}-(L^{-1})_{12}\\ L_{21}&L_{22}\end{pmatrix}\approx\begin{pmatrix}L_{11}&(1+\frac{1}{C_{0}^{L}})L_{12}\\ L_{21}&L_{22}\end{pmatrix},\quad\det(L^{(1)})\approx C_{0}^{(1)L}+C_{1}^{(1)L}\delta k,$
    $\det(h_{0}^{(1)})$ $=$ $\det(R^{(1)})\det(L^{(1)})\approx C_{0}^{(1)R}C_{0}^{(1)L}+(C_{0}^{(1)R}C_{1}^{(1)L}+C_{0}^{(1)L}C_{1}^{(1)R})\color[rgb]{0,0,1}\delta k.$ (S21)
@@
    $[R^{(n)}]_{21}$ $=$ $R_{21}-[(R^{(n-1)})^{-1}]_{21}=R_{21}+\frac{[R^{(n-1)}]_{21}}{\det(R^{(n-1)})}\sim R_{21},\quad\text{ and similarly, }\quad[L^{(n)}]_{12}\sim L_{12},$ (S22)
    which ensures that the expansion of $\det(h_{0}^{(n)})=\det(R^{(n)})\det(L^{(n)})\sim\det(h_{0})$ always remains in the form of $C_{0}^{(n)}+C_{1}^{(n)}\delta k$. Therefore, in this case, we can obtain the linear dispersion of $\det(H_{\text{y-OBC}})$ as:
    $\det(H_{\text{y-OBC}})$ $=$ $\det(h_{0})\det(h_{0}^{(1)})\cdots\det(h_{0}^{(L_{y}-1)})$ (S23)
    $\approx$ $(C_{0}+C_{1}\delta k)*(C_{0}^{(1)}+C_{1}^{(1)}\delta k)\cdots(C_{0}^{(L_{y}-1)}+C_{1}^{(L_{y}-1)}\delta k)\sim\color[rgb]{0,0,1}\delta k,$
    where the constant term $C_{0}C_{0}^{(1)}\cdots C_{0}^{(L_{y}-1)}$ goes to zero due to the existence of zero energy modes as shown in Fig. S1(a).
The above approach hinges on the observation that the two types of dispersion are distinguished by whether $\text{det}{(R)}=0$. It works for models where most matrix elements in $h_{+}^{y}$ and $h_{-}^{y}$ are 00, which effectively simplifies the calculation of $h_{0}^{(n)}=h_{0}-h_{+}[h_{0}^{(n-1)}]^{-1}h_{-}$ in Eq. S11. In more general models with hoppings beyond nearest neighbors, Eq. S11 would need to be extended to handle multiple matrix diagonals, and this iterative approach could become far more complicated.
II II. Detailed analysis of the topological flat band model

### v4

@@ II.1.3 Deriving two types of dispersion relations (linear and square-root) for topological edge bands
For the general form of the Hamiltonian $H(k,k_{y})$ as defined in Eq (24), when considering OBCs in the $y$-direction with size $L_{y}$, we obtain:
$H_{\text{y-OBC}}(k)=\begin{pmatrix}h_{0}^{y}&h_{+}^{y}&\cdots&0&0\\ h_{-}^{y}&h_{0}^{y}&h_{+}^{y}&\cdots&0\\ 0&h_{-}^{y}&h_{0}^{y}&\cdots&0\\ \vdots&\vdots&\ddots&\ddots&\vdots\\ 0&0&\cdots&h_{-}^{y}&h_{0}^{y}\end{pmatrix}_{L_{y}\times L_{y}},$ (25)
with
$h_{+}^{y}=\begin{pmatrix}0&0&0&0\\ 0&0&0&1\\ 1&0&0&0\\ 0&0&0&0\end{pmatrix},h_{-}^{y}=\begin{pmatrix}0&0&1&0\\ 0&0&0&0\\ 0&0&0&0\\ 0&1&0&0\end{pmatrix},\text{ and }h_{0}^{y}(k)=\begin{pmatrix}0&R\\ L&0\end{pmatrix},$ (26)
@@
$R=\begin{pmatrix}f+\delta+Z&2Z-i(Z+\sin(k+k_{0})\\ i(Z-\sin(k+k_{0}))&f+\delta-Z\end{pmatrix},$
$L=\begin{pmatrix}f-\delta+Z&-i(Z-\sin(k+k_{0})\\ 2Z+i(Z+\sin(k+k_{0}))&f-\delta-Z\end{pmatrix},$ (27)
and $f=\cos(k+k_{0})+\cos k_{0}-M$. For this $H_{\text{y-OBC}}(k)$, the topological edge bands exhibit two different types of dispersion relations around the EP: (1) a linear dispersion $E_{e}(\delta k)\sim\delta k$, as shown in Fig. 3(c), and (2) square-root dispersion $E_{e}(\delta k)\sim\sqrt{\delta k}$, as depicted in Fig. 4(a).
To explain why the dispersion can exhibit two qualitatively different behavior (i.e., becomes linear when we set $Z=1,\delta=M-Z$ in Eq (24)), we expand the momentum $k$ around the EP as: $k\rightarrow 0+\delta k$. We shall prove that:
$\det[H_{\text{y-OBC}}(\delta k)]\sim\begin{cases}\delta k^{2},&\text{with }Z=1,\delta=M-Z\\ \delta k,&\text{other parameters}\end{cases}.$ (28)
This scaling behavior is directly linked to the dispersion of the topological bands, which are the only bands that become gapless in the spectrum. To elaborate, we know that the determinant is equal to the product of all eigenenergies as:
$\det[H_{\text{y-OBC}}(k)]=E_{e_{1}}(k)E_{e_{2}}(k)\prod_{n}E_{n}(k),$ (29)
@@
To derive Eq (28), we first expand $H_{\text{y-OBC}}$ around $k=0$. Since $k$ appears only in the diagonal element $h_{0}^{y}$, we need to expand the $R$ and $L$ matrices [Eq (27)]. By substituting $\sin k_{0}=Z$ into Eq (27), we obtain
$\sin(k_{0}+\delta k)\approx Z+\cos k_{0}\delta k-\frac{Z}{2}\delta k^{2},$
$\cos(k_{0}+\delta k)\approx\cos k_{0}-Z\delta k-\frac{\cos k_{0}}{2}\delta k^{2},$ (30)
@@
$R(\delta k)$ $=$ $\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}$ (31)
$=$ $\begin{pmatrix}f_{0}+\delta+Z&2Z(1-i)\\ 0&f_{0}+\delta-Z\end{pmatrix}+\begin{pmatrix}-Z&-i\cos k_{0}\\ -i\cos k_{0}&-Z\end{pmatrix}\delta k$
$+\begin{pmatrix}-\frac{1}{2}\cos k_{0}&i\frac{Z}{2}\\ i\frac{Z}{2}&-\frac{1}{2}\cos k_{0}\end{pmatrix}\delta k^{2}+\mathcal{O}(\delta k^{3}),$
@@
$L(\delta k)$ $=$ $\begin{pmatrix}L_{11}&L_{12}\\ L_{21}&L_{22}\end{pmatrix}$ (32)
$=$ $\begin{pmatrix}f_{0}-\delta+Z&0\\ 2Z(1+i)&f_{0}-\delta-Z\end{pmatrix}+\begin{pmatrix}-Z&i\cos k_{0}\\ i\cos k_{0}&-Z\end{pmatrix}\delta k$
$+\begin{pmatrix}-\frac{1}{2}\cos k_{0}&-i\frac{Z}{2}\\ -i\frac{Z}{2}&-\frac{1}{2}\cos k_{0}\end{pmatrix}\delta k^{2}+\mathcal{O}(\delta k^{3}).$
@@
$\det$ $\begin{pmatrix}h_{0}&h_{+}&\cdots&0\\ h_{-}&h_{0}&h_{+}&\vdots\\ 0&h_{-}&\ddots&h_{+}\\ 0&\cdots&h_{-}&h_{0}\end{pmatrix}_{L_{y}}=\det(h_{0})\det\begin{pmatrix}h_{0}&h_{+}&\cdots&0\\ h_{-}&h_{0}&h_{+}&\vdots\\ 0&h_{-}&\ddots&h_{+}\\ 0&\cdots&h_{-}&\color[rgb]{0,0,1}h_{0}^{(1)}\end{pmatrix}_{L_{y}-1}$ (33)
$=$ $\det(h_{0})\det(h_{0}^{(1)})\det\begin{pmatrix}h_{0}&h_{+}&\cdots&0\\ h_{-}&h_{0}&h_{+}&\vdots\\ 0&h_{-}&\ddots&h_{+}\\ 0&\cdots&h_{-}&\color[rgb]{0,0,1}h_{0}^{(2)}\end{pmatrix}_{L_{y}-2}$
@@
in which $h_{0}^{(1)}=h_{0}-h_{+}h_{0}^{-1}h_{-}=\begin{pmatrix}0&R^{(1)}\\ L^{(1)}&0\end{pmatrix}$, with
$R^{(1)}$ $=$ $R-\begin{pmatrix}0&0\\ (R^{-1})_{21}&0\end{pmatrix},\quad L^{(1)}=L-\begin{pmatrix}0&(L^{-1})_{12}\\ 0&0\end{pmatrix}$ (34)
and $h_{0}^{(2)}=h_{0}-h_{+}[h_{0}^{(1)}]^{-1}h_{-}=\begin{pmatrix}0&R^{(2)}\\ L^{(2)}&0\end{pmatrix}$, with
$R^{(2)}=R-\begin{pmatrix}0&0\\ [(R^{(1)})^{-1}]_{21}&0\end{pmatrix},L^{(2)}=L-\begin{pmatrix}0&[(L^{(1)})^{-1}]_{12}\\ 0&0\end{pmatrix},$ (35)
and so forth $h_{0}^{(n)}=h_{0}-h_{+}[h_{0}^{(n-1)}]^{-1}h_{-}$. Now, to calculate Eq (33), we need to determine how $\text{det}(h_{0}^{(n)})$ varies with $\delta k$.
@@
    $R(\delta k)=\begin{pmatrix}-\delta k&2(1-i)+\frac{i}{2}\delta k^{2}\\ \frac{i}{2}\delta k^{2}&-2-\delta k\end{pmatrix}+\mathcal{O}(\delta k^{3}),$
    $L(\delta k)=\begin{pmatrix}2(1-M)-\delta k&-\frac{i}{2}\delta k^{2}\\ 2(1+i)-\frac{i}{2}\delta k^{2}&-2M\end{pmatrix}+\mathcal{O}(\delta k^{3}),$ (36)
@@
    $\det(h_{0})=\det(R)\det(L)\approx-8M(1-M)\delta k.$ (37)
    Importantly, the fact that $\det(R)$ is proportional to $2\delta k$ and vanishes as $\delta k\rightarrow 0$ will turn out to be crucial. According to Eq (35), for the first iteration, $\det(h_{0}^{(1)})$, can be calculated using $R^{(1)}$ and $L^{(1)}$
    $R^{(1)}$ $=$ $\begin{pmatrix}R_{11}&R_{12}\\ R_{21}-(R^{-1})_{21}&R_{22}\end{pmatrix}=\begin{pmatrix}R_{11}&R_{12}\\ (1+\frac{1}{\det(R)})R_{21}&R_{22}\end{pmatrix}$
    $\approx$ $\begin{pmatrix}-\delta k&2(1-i)+\frac{i}{2}\delta k^{2}\\ \frac{i}{4}\delta k+\frac{i}{2}\delta k^{2}&-2-\delta k\end{pmatrix}+\mathcal{O}(\delta k^{3}),$
    $L^{(1)}$ $=$ $\begin{pmatrix}L_{11}&L_{12}-(L^{-1})_{12}\\ L_{21}&L_{22}\end{pmatrix}=\begin{pmatrix}L_{11}&\left(1-\frac{1}{4M(1-M)}\right)L_{12}\\ L_{21}&L_{22}\end{pmatrix}$ (38)
    $\approx$ $\begin{pmatrix}2(1-M)-\delta k&-\frac{i}{2}\left(1-\frac{1}{4M(1-M)}\right)\delta k^{2}\\ 2(1+i)-\frac{i}{2}\delta k^{2}&-2M\end{pmatrix}$
    $+\mathcal{O}(\delta k^{3}),$
@@
    $\det(h_{0}^{(1)})$ $=$ $\det(R^{(1)})\det(L^{(1)})$ (39)
    $\approx$ $-4M(1-M)\left(2+\frac{1+i}{2}\right)\delta k.$
@@
$[R^{(2)}]_{21}$ $=$ $R_{21}-[(R^{(1)})^{-1}]_{21}=R_{21}+\frac{R^{(1)}_{21}}{\det(R^{(1)})}$
$\approx$ $\frac{i}{10+2i}=\text{Const},$
$\det(R^{(2)})$$\approx$ $-\frac{1+i}{5+i}=\text{Const},$
$[R^{(n)}]_{21}$ $=$ $R_{21}-[(R^{(n-1)})^{-1}]_{21}=R_{21}+\frac{R^{(n-1)}_{21}}{\det(R^{(n-1)})}$
$=$ $\frac{i}{2}\delta k^{2}+\frac{\text{Const}}{\text{Const}}\approx\text{Const},$
$\det(R^{(n)})|_{n\geq 2}$$\approx$ $2(1-i)*[R^{(n)}]_{21}=\text{Const},$ (40)
@@
$[L^{(2)}]_{12}$ $=$ $L_{12}-[(L^{(1)})^{-1}]_{12}=-\frac{i\delta k^{2}}{2}+\frac{L_{12}^{(1)}}{\det(L^{(1)})}$
$\approx$ $\frac{i\delta k^{2}}{2}\left(\frac{1-\frac{1}{4M(1-M)}}{4M(1-M)}-1\right),$
$\det(L^{(2)})$$\approx$ $-4M(1-M)=\text{Const},$
$[L^{(n)}]_{12}$ $=$ $L_{12}+\frac{L_{12}^{(n-1)}}{\det(L^{(n-1)})}\sim\delta k^{2},$
$\det(L^{(n)})|_{n\geq 2}$$\approx$ $-2(1-M)*2M=-4M(1-M)$ (41)
[…]
@@
$\det(H_{\text{y-OBC}})=\det(h_{0})\det(h_{0}^{(1)})\cdots\det(h_{0}^{(L_{y}-1)})\sim\delta k^{2}.$ (42)
  • For most other generic parameter combinations, such as the case shown in Fig. 4(a), the constant term in $\det(R)$ does not vanish i.e., $\det(R)\rightarrow\text{Const}$ as $\delta k\rightarrow 0$. Below we show how that leads to a qualitatively different dispersion. We have
    $\det(R)$$\approx$ $(f_{0}+\delta)^{2}-Z^{2}-2Z\big{(}f_{0}+\delta-(1+i)\cos k_{0}\big{)}\delta k$
    $=$ $C_{0}^{R}+C_{1}^{R}\delta k,$
    $\det(L)$$\approx$ $(f_{0}-\delta)^{2}-Z^{2}-2Z\big{(}f_{0}-\delta+(1-i)\cos k_{0}\big{)}\delta k$
    $=$ $C_{0}^{L}+C_{1}^{L}\delta k,$
    $\det(h_{0})$ $=$ $\det(R)\det(L)$ (43)
    $\approx$ $C_{0}^{R}C_{0}^{L}+(C_{0}^{R}C_{1}^{L}+C_{0}^{L}C_{1}^{R})\delta k.$
@@
    $R^{(1)}$ $=$ $\begin{pmatrix}R_{11}&R_{12}\\ R_{21}-(R^{-1})_{21}&R_{22}\end{pmatrix}\approx\begin{pmatrix}R_{11}&R_{12}\\ (1+\frac{1}{C_{0}^{R}})R_{21}&R_{22}\end{pmatrix},$
    $\det(R^{(1)})$$\approx$ $C_{0}^{(1)R}+C_{1}^{(1)R}\delta k,$
    $L^{(1)}$ $=$ $\begin{pmatrix}L_{11}&L_{12}-(L^{-1})_{12}\\ L_{21}&L_{22}\end{pmatrix}\approx\begin{pmatrix}L_{11}&(1+\frac{1}{C_{0}^{L}})L_{12}\\ L_{21}&L_{22}\end{pmatrix},$
    $\det(L^{(1)})$$\approx$ $C_{0}^{(1)L}+C_{1}^{(1)L}\delta k,$
    $\det(h_{0}^{(1)})$ $=$ $\det(R^{(1)})\det(L^{(1)})$ (44)
    $\approx$ $C_{0}^{(1)R}C_{0}^{(1)L}+(C_{0}^{(1)R}C_{1}^{(1)L}+C_{0}^{(1)L}C_{1}^{(1)R})\delta k.$
@@
    $[R^{(n)}]_{21}$ $=$ $R_{21}-[(R^{(n-1)})^{-1}]_{21}=R_{21}+\frac{[R^{(n-1)}]_{21}}{\det(R^{(n-1)})}$ (45)
    $\sim$ $R_{21},$
    and similarly, $[L^{(n)}]_{12}\sim L_{12}$, which ensures that the expansion of $\det(h_{0}^{(n)})=\det(R^{(n)})\det(L^{(n)})\sim\det(h_{0})$ always remains in the form of $C_{0}^{(n)}+C_{1}^{(n)}\delta k$. Therefore, in this case, we can obtain the linear dispersion of $\det(H_{\text{y-OBC}})$ as:
    $\det$ $(H_{\text{y-OBC}})=\det(h_{0})\det(h_{0}^{(1)})\cdots\det(h_{0}^{(L_{y}-1)})$ (46)
    $\approx$ $(C_{0}+C_{1}\delta k)*(C_{0}^{(1)}+C_{1}^{(1)}\delta k)\cdots(C_{0}^{(L_{y}-1)}+C_{1}^{(L_{y}-1)}\delta k)$
    $\sim$ $\delta k,$
    where the constant term $C_{0}C_{0}^{(1)}\cdots C_{0}^{(L_{y}-1)}$ goes to zero due to the existence of zero energy modes as shown in Fig. 4(a).
The above approach hinges on the observation that the two types of dispersion are distinguished by whether $\text{det}{(R)}=0$. It works for models where most matrix elements in $h_{+}^{y}$ and $h_{-}^{y}$ are 00, which effectively simplifies the calculation of $h_{0}^{(n)}=h_{0}-h_{+}[h_{0}^{(n-1)}]^{-1}h_{-}$ in Eq (33). In more general models with hoppings beyond nearest neighbors, Eq (33) would need to be extended to handle multiple matrix diagonals, and this iterative approach could become far more complicated.
II.2 Detailed analysis of the topological flat band model